# A Survey of Frontiers in LLM Reasoning: Inference Scaling, Learning to Reason, and Agentic Systems

### v1

@@ Abstract
Figure 1: Growth trend in LLM reasoning. We show the cumulative number (in thousands) of papers published from 2022 to February 2025, based on Semantic Scholar keyword search. Research on regimes and architectures has accelerated notably since the introduction of Chain-of-Thought (CoT) in 2022.
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated a range of emerging abilities, such as in-context learning (Dong et al., 2024), and role playing (Shanahan et al., 2023b) as they scale, with reasoning becoming one of the most critical capabilities. As shown in Figure 1, this area has rapidly gained research attention, often referred to as LLM reasoning or reasoning language model (RLM) (Besta et al., 2025). The increasing focus on this topic is understandable, as reasoning capability is: (i) Challenging, requiring multi-step processing beyond the token-by-token generative nature of auto-regressive LLMs; (ii) Fundamental, as it is a core aspect of intelligence, particularly in planning and strategic decision-making; and, most importantly, (iii) Promising, as recent advances in LLMs hint at a viable path forward. Given these factors, reasoning is widely regarded as a prerequisite for more advanced AI systems approaching Artificial General Intelligence (AGI), beyond the conventional AI that aims to closely follow instruction (Duenas & Ruiz, 2024).
@@ 1.2 Reasoning System Architecture
To distinguish agentic systems from standalone LLMs, we adopt the perspective of Kapoor et al. (2024), framing agentic behavior as a spectrum. We categorize these systems into two families: single-agent and multi-agent. In single-agent systems, a single LLM interacts with tools in its environment to refine reasoning, actions, and perceptions. These tools include external knowledge bases (Hammane et al., 2024; Sun et al., 2023), verifiers (Wan et al., 2024c; Guan et al., 2025), and practical applications like code interpreters, calendars, and maps (Yu et al., 2023b; Lu et al., 2024a). By leveraging these resources, the LLM iteratively enhances its decision-making and problem-solving capabilities. Recent milestones in single-agent systems, such as Grok 3 Deep Search (02/2025) and OpenAI Deep Research (02/2025), demonstrate how agents interact with the web to significantly improve reasoning, perform tasks like information retrieval, use code interpreters for calculations, and aggregate data from multiple sources.
@@ 1.4 Goal and Structure of the Survey
LLM §5.1, draw=vermillion, fill=vermillion!10, text width=2.2cm, align=left [Training from Trajectories: Huang et al. (2024d); Wang et al. (2023d), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Trajectory Collection: Zelikman et al. (2022); Yuan et al. (2023); Qin et al. (2024b), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Prompt Construction: Xu et al. (2024a); Yu et al. (2024c); Li et al. (2024a), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] ] ] [Learning
@@
Verifiers §4.2, draw=reddishpurple, fill=reddishpurple!10, text width=2.2cm, align=left [Generative Verifiers: Yao et al. (2023b); Zhang et al. (2024f); Mahan et al. (2024), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] [Process Reward Models: Wang et al. (2024g); Jiao et al. (2024a); Setlur et al. (2024b), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] [Outcome Reward Models: Liu et al. (2024a), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] ] [Learning of
@@ 1.5 Comparison to Related Surveys
Reasoning in LLMs has long been a fundamental challenge in the field. Huang & Chang (2023) provide a comprehensive overview of the evolution of informal deductive reasoning, tracing its development prior to the rise of LLM agents and Reasoning Language Models (RLMs). Qiao et al. (2023b) offer a detailed summary of advancements in LLM reasoning, with a particular emphasis on prompting techniques. In contrast, Yu et al. (2024a) distinguish their work by focusing on establishing a formal definition and taxonomy of natural language reasoning, rooted in both philosophical foundations and real-world applications.
@@
Improvements in LLM reasoning are closely tied to advancements in a variety of techniques. Dong et al. (2024) present a comprehensive survey on in-context learning (ICL), while Zhou et al. (2024b) explore the interpretation and analysis of ICL from both theoretical and empirical perspectives. Recent studies suggest that enhancements in reasoning are often linked to inference scaling. Dong et al. (2024) provide an extensive review of inference-time self-improvement, and Welleck et al. (2024) offer a survey focused on three key themes: token-level generation algorithms, meta-generation algorithms, and efficient generation. Following the release of Reasoning Language Models (RLMs) such as OpenAI’s o1 and DeepSeek’s R1, there has been a significant increase in research dedicated to learning-to-reason approaches. Zeng et al. (2024) and Xu et al. (2025c) provide thorough surveys on these emerging developments.
@@
Research on LLM reasoning has predominantly centered on logical and mathematical reasoning. Liu et al. (2025a) offer a comprehensive survey of logical reasoning in LLMs, delving into its theoretical foundations and associated benchmarks. In their position paper, Yang et al. (2024d) underscore the pivotal role of formal mathematical reasoning, showcasing its superiority over traditional NLP-based methods in generating verifiable proofs and automated feedback. Their work outlines progress in theorem proving and auto-formalization while identifying key challenges that remain. Pezeshkpour et al. (2024a) contribute a positioning paper on the reasoning capacity of multi-agent systems, defining it as the ability to effectively gather input data, process information, and produce accurate outputs for specific tasks under constraints.
@@
A concurrent work by Besta et al. (2025) introduces a comprehensive and modular framework for RLMs that systematically organizes key components such as reasoning structures, strategies, benchmarks and learning algorithms. However, their work does not delve into agentic and multi-agent LLM systems. [^1] To avoid redundancy with existing literature, we do not include an analysis of reasoning benchmarks in this survey. For a detailed discussion of benchmarks, we direct readers to Xu et al. (2025c); Besta et al. (2025).
@@
Reasoning is a critical capability in agentic systems (Pezeshkpour et al., 2024b; Masterman et al., 2024). While numerous reviews focus on agent systems (Xi et al., 2023; Kapoor et al., 2024), discussions on reasoning within these systems remain limited. This survey provides a comprehensive overview of major milestones in LLM reasoning research, emphasizing two key dimensions: (1) the evolution of learning schemes—from inference scaling to learning-to-reason approaches—and (2) architectural advancements—from single LLMs to multi-agent systems. These dimensions summarize recent progress and lay the groundwork for future reasoning LLMs and agentic systems. We unify techniques under input and output perspectives, clarifying what must be customized or designed when building reasoning systems. Additionally, we detail essential techniques, including a comparison of the latest learning algorithms (e.g., RL) and an in-depth discussion of refiners and verifiers, which are critical for facilitating reasoning. Given these contributions, our survey is timely, offering AI researchers up-to-date insights into the field. We anticipate further research along these dimensions, such as agent-human regimes (Liang et al., 2024) and automated workflow design architectures (Hu et al., 2025; Zhang et al., 2024c; Zhou et al., 2025).
@@ 2.4 Reasoning Regimes
Figure 6: Inference-time and training-time regimes of a reasoning system. We use tree search as an example to illustrate the inference scaling and trajectories collection. Given a query, inference scaling relies on extensive inference computation to improve the reasoner’s distribution. Specifically, it generates multiple candidate reasoning steps at each layer and sselects the best solution to proceed (e.g., by using an external verifier or ensembling). In contrast, learning to reason focuses on collecting trajectories and training from the collected data with minimal inference-time computation. It takes all trajectories in the process and labels them with preferences. The preference data can then be used to train the reasoner.
@@ 3.1.1 Constructing Reasoning Provoking Prompts
One simple yet effective way to improve the reasoning abilities of LLMs is to instruct them to explicitly conduct reasoning, often with the help of few-shot exemplars.
@@
Humans can address new problems by drawing analogy from relevant past experience (Holyoak, 2012). Inspired by this, Yasunaga et al. (2024) propose analogical prompting to guide LLMs to self-generate exemplars or knowledge relevant to the given problem as few-shot demonstrations for reasoning, outperforming hand-crafted or retrieved examples. For example, LLMs are prompted to generate a problem on calculating a third-order determinant before solving the given fourth-order determinant. Similarly, Chen et al. (2023d); Yang et al. (2023a); Luo et al. (2024a) highlight the effectiveness of self-generated relevant exemplars. Qin et al. (2024a) further systematically assess the capability of LLMs to perform analogical reasoning, revealing that LLMs cannot always perform analogical reasoning. They identify the quality of the self-generated exemplars as the limiting factor, rather than exemplar relevance.
Conventionally, a fixed set of few-shot demonstrations is applied to all queries, which can be suboptimal, especially when queries vary significantly. An alternative approach is to retrieve demonstrations tailored to the current query. Research has shown that retrieval-based demonstration selection significantly improves task performance. The main goals for selecting demonstrations are similarity (Rubin et al., 2022; Agrawal et al., 2023; Li et al., 2023d; Ye et al., 2023a) and diversity (Levy et al., 2023; He et al., 2023; Kim et al., 2024a). Various retrieval strategies have been proposed for selecting $k$ demonstrations, including top-$k$ similarity-based retrieval (Liu et al., 2022; Li et al., 2023d), clustering-based retrieval (Luo et al., 2023c; Wang et al., 2024i), and iterative retrieval (Khattab et al., 2022; Levy et al., 2023; Wang et al., 2024e). These methods enable adaptive and effective demonstration selection, enhancing the model’s reasoning and generalization across diverse queries.
@@ 3.1.2 Optimizing Reasoning Output with Search and Planning
Beyond subtasks, researchers emphasize the importance of robust reasoning structures such as hierarchical and decision-making processes that capture the underlying mechanisms involved in problem-solving. Zhou et al. (2024a) introduce Self-Discover, a framework that enables models to self-identify reasoning structures for any task using a seed set of general reasoning skill modules. Building on this, Aswani et al. (2024) propose Auto-Evolve, which dynamically adapts reasoning modules to accommodate more diverse problems. In addition to designing better reasoning steps, several studies address the need to correct intermediate steps. For example, Deng et al. (2024a); Yan et al. (2024) and Wu et al. (2024b) propose methods to refine intermediate outputs. Notably, Zhang et al. (2024i) observe that smaller models ($\leq$ 13B parameters) in particular need stronger models acting as verifiers to validate and correct intermediate steps.
@@ 4.1.2 Reinforcement Learning for Reasoning
At each time step $t$, for a given state $s_{t}\in\mathcal{S}$, the agent selects an action $a_{t}$ and transitions to a new state $s_{t+1}$, receiving a reward $R(s_{t},a_{t})$ from the environment. The set of available actions at state $s_{t}$ may be restricted to a subset of of $\mathcal{A}$, denoted $\mathcal{A}_{s_{t}}$ (i.e., $a_{t}\in\mathcal{A}_{s_{t}}$). A key assumption of MDPs is that the current state $s_{t}$ fully encapsulates all relevant information about the environment. This means the next state $s_{t+1}$ depends solely on the current state $s_{t}\in\mathcal{S}$ and the chosen action $a_{t}\in\mathcal{A}_{s_{t}}$. As such, the state transition is agnostic to the history or previous states and actions. Within this MDP framework, the goal of RL is to learn a policy model that selects optimal actions by maximizing the expected cumulative rewards (Eq. 1).
In training large language models (LLMs) for reasoning, approaches vary based on how they define states and actions. These methods can be grouped into three categories as below (summarized in Table 5):
@@
As one of the primary variants of policy gradient methods, PPO has remained a popular and widely used RL algorithm (Schulman et al., 2017). To train the policy $\pi_{\theta}$, PPO utilizes two additional models: the reference model $\pi_{\theta_{\text{ref}}}$, which represents the initial state of the policy, and the value model $V$, which estimates the state value $V(s_{t})$. PPO begins by sampling a state-action trajectory $\tau$ with consecutive state-action pairs $s_{t+1}\sim(s_{t},a_{t})$, then collects the respective intermediate or process reward (if available) and final (outcome) reward. Then, it computes the advantage $A(s_{t},a_{t})$ of each action $a_{t}$ given the current state $s_{t}$, which is defined as the relative strength of that specific action $a_{t}$ compared to the average actions that could have been taken from $s_{t}$. The advantage is formulated as
@@
where $Q(s_{t},a_{t})$ represents the expected cumulative total reward that the policy is expected to obtain if it takes action $a_{t}$ from $s_{t}$, while $V(s_{t})$ denotes the expected total rewards obtainable from state $s_{t}$, known as the state value. The state value is equivalent to the expected value of $Q(s_{t},a^{\prime}_{t})$ marginalized over all possible actions available from $s_{t}$. If $A(s_{t},a_{t})>0$, the action $a_{t}$ is encouraged, conversely, if $A(s_{t},a_{t})<0$, the action $a_{t}$ is discouraged. After computing the advantages, PPO optimizes the policy $\pi_{\theta}$ according to the following loss function.
@@
where $t\in[0,T]$ is a time step within trajectory $\tau$, $\pi_{\theta_{o}}$ is the fixed policy from previous episode or iteration, and $P$ is the transition distribution. The $clip$ function, applied to the probability ratio $\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{o}}(a_{t}|s_{t})}$, ensures that the policy does not deviate too drastically or rapidly from its previous version. This also help prevent catastrophic failure or suboptimal local solutions. Additionally, a KL divergence term $\mathcal{D_{\text{KL}}}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})$ is often incorporated into the loss function to constrain exploration during the later stages of training. Throughout the training process, both the policy $\pi_{\theta}$ and value model $V$ are iteratively updated.
@@
$L_{REINFORCE}(\theta)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}[(R(y,x)-b)\nabla_{\pi_{\theta}}\log\pi_{\theta}(y|x)]$ (5)
@@
$L_{RLOO}(\theta)=\frac{1}{k}\sum_{i=1}^{k}[R(y_{i},x)-\frac{1}{k-1}\sum_{j\neq i}R(y_{j},x)]\nabla_{\pi_{\theta}}\log\pi_{\theta}(y_{i}|x)$ (6)
where $k$ denotes the number of Monte Carlo samples. Unlike PPO, these algorithms do not rely on a parameterized value function and instead depend solely on observed rewards. These methods share similarities with approaches such as Group-Relative Policy Optimization (GRPO) (Ramesh et al., 2024) and VinePPO (Kazemnejad et al., 2024), which will be discussed in detail below.

### v2

@@ Abstract
Figure 1: The LLM reasoning surge. We show the cumulative number (in thousands) of papers published from 2022 to 2/2025, based on Semantic Scholar keyword search. Research on reasoning regimes and agent architectures has accelerated notably since the introduction of Chain-of-Thought (CoT) in 2022. This growth is further influenced by other major developments, such as the release of ChatGPT (Ouyang et al., 2022) in 9/2022, and popularity of in-context learning (Brown et al., 2020) as an inference-time optimization method.
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated a range of emerging abilities, such as in-context learning (Dong et al., 2024), role playing (Shanahan et al., 2023b) and domain adaptation (Ke et al., 2023; 2025a; Ke & Liu, 2023) as they scale, with reasoning becoming one of the most critical capabilities. As shown in Figure 1, this area has rapidly gained research attention, often referred to as LLM reasoning or reasoning language model (RLM) (Besta et al., 2025). The increasing focus on this topic is understandable, as reasoning capability is: (i) Challenging, requiring multi-step processing beyond the token-by-token generative nature of auto-regressive LLMs; (ii) Fundamental, as it is a core aspect of intelligence, particularly in planning and strategic decision-making; and, most importantly, (iii) Promising, as recent advances in LLMs hint at a viable path forward. Given these factors, reasoning is widely regarded as a prerequisite for more advanced AI systems approaching Artificial General Intelligence (AGI), beyond the conventional AI that aims to closely follow instruction (Duenas & Ruiz, 2024).
@@ 1.2 Reasoning System Architecture
To distinguish agentic systems from standalone LLMs, we adopt the perspective of Kapoor et al. (2024), framing agentic behavior as a spectrum. We categorize these systems into two families: single-agent and multi-agent. In single-agent systems, a single LLM interacts with tools in its environment to refine reasoning, actions, and perceptions. These tools include external knowledge bases (Ke et al., 2024; Hammane et al., 2024; Sun et al., 2023), verifiers (Wan et al., 2024c; Guan et al., 2025), and practical applications like code interpreters, calendars, and maps (Yu et al., 2023b; Lu et al., 2024a). By leveraging these resources, the LLM iteratively enhances its decision-making and problem-solving capabilities. Recent milestones in single-agent systems, such as Grok 3 Deep Search (02/2025) and OpenAI Deep Research (02/2025), demonstrate how agents interact with the web to significantly improve reasoning, perform tasks like information retrieval, use code interpreters for calculations, and aggregate data from multiple sources.
@@ 1.4 Goal and Structure of the Survey
LLM §5.1, draw=vermillion, fill=vermillion!10, text width=2.2cm, align=left [Training from Trajectories: Huang et al. (2024d); Wang et al. (2023d), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Trajectory Collection: Zelikman et al. (2022); Yuan et al. (2023); Qin et al. (2024), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] [Prompt Construction: Xu et al. (2024a); Yu et al. (2024c); Li et al. (2024a), draw=vermillion, fill=vermillion!5, text width=11.2cm, align=left] ] ] [Learning
@@
Verifiers §4.2, draw=reddishpurple, fill=reddishpurple!10, text width=2.2cm, align=left [Generative Verifiers: Yao et al. (2023c); Zhang et al. (2024f); Mahan et al. (2024), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] [Process Reward Models: Wang et al. (2024g); Jiao et al. (2024a); Setlur et al. (2024b), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] [Outcome Reward Models: Liu et al. (2024a), draw=reddishpurple, fill=reddishpurple!5, text width=11.2cm, align=left] ] [Learning of
@@ 1.5 Comparison to Related Surveys
Reasoning in LLMs has long been a fundamental challenge in the field. Earlier works, such as Huang & Chang (2023), provide a comprehensive overview of the evolution of informal deductive reasoning covering developments prior to the emergence of LLM agents and Reasoning Language Models (RLMs). Our work extends this discussion by focusing on LLM agents and RLMs. Qiao et al. (2023b) offer a detailed summary of advancements in LLM reasoning, with a particular emphasis on prompting techniques. In contrast, we offer a broader range of regimes (from inference to training) and architectures (from standalone LLM to multi-agent systems). Readers interested in a formal definition and taxonomy of natural language reasoning—grounded in philosophical foundations—may refer to Yu et al. (2024a), which focuses specifically on this direction and is complementary to our scope.
@@
Improvements in LLM reasoning are closely tied to advancements in a variety of techniques. Dong et al. (2024) present a comprehensive survey on in-context learning (ICL), while Zhou et al. (2024c) explore the interpretation and analysis of ICL from both theoretical and empirical perspectives. In contrast, our work organizes ICL techniques under different regimes—standalone LLMs, single-agent, and multi-agent systems—highlighting how these techniques evolve and interact within each setting. Recent studies suggest that enhancements in reasoning are often linked to inference scaling. Dong et al. (2024) provide an extensive review of inference-time self-improvement, and Welleck et al. (2024) offer a survey focused on three key themes: token-level generation algorithms, meta-generation algorithms, and efficient generation. Following the release of Reasoning Language Models (RLMs) such as OpenAI’s o1 and DeepSeek’s R1, there has been a significant increase in research dedicated to learning-to-reason approaches. Zeng et al. (2024) and Xu et al. (2025d) provide thorough surveys on these emerging developments. However, these surveys primarily focus on LLMs, and do not address agentic or multi-agent reasoning settings in depth.
@@
Research on LLM reasoning has predominantly centered on logical and mathematical reasoning. Liu et al. (2025a) offer a comprehensive survey of logical reasoning in LLMs, delving into its theoretical foundations and associated benchmarks. In their position paper, Yang et al. (2024d) underscore the pivotal role of formal mathematical reasoning, showcasing its superiority over traditional NLP-based methods in generating verifiable proofs and automated feedback. Their work outlines progress in theorem proving and auto-formalization while identifying key challenges that remain. While we cover domain-specific reasoning in Section 6.1.3, we refer readers to Liu et al. (2025a) and Yang et al. (2024d) for a more in-depth treatment of these topics.
@@
Reasoning is a critical capability in agentic systems (Pezeshkpour et al., 2024; Masterman et al., 2024). While numerous reviews focus on agent systems (Xi et al., 2023; Kapoor et al., 2024), discussions on reasoning within these systems remain limited. A concurrent work by Besta et al. (2025) introduces a comprehensive and modular framework for RLMs that systematically organizes key components such as reasoning structures, strategies, benchmarks and learning algorithms. However, their work does not delve into agentic and multi-agent LLM systems. [^1] To avoid redundancy with existing literature, we do not include an analysis of reasoning benchmarks in this survey. For a detailed discussion of benchmarks, we direct readers to Xu et al. (2025d); Besta et al. (2025).
@@
This survey provides a comprehensive overview of major milestones in LLM reasoning research, emphasizing two key dimensions: (1) the evolution of learning schemes—from inference scaling to learning-to-reason approaches—and (2) architectural advancements—from single LLMs to multi-agent systems. These dimensions summarize recent progress and lay the groundwork for future reasoning LLMs and agentic systems. We unify techniques under input and output perspectives, clarifying what must be customized or designed when building reasoning systems. Additionally, we detail essential techniques, including a comparison of the latest learning algorithms (e.g., RL) and an in-depth discussion of refiners and verifiers, which are critical for facilitating reasoning. Given these contributions, our survey is timely, offering AI researchers up-to-date insights into the field. We anticipate further research along these dimensions, such as agent-human regimes (Liang et al., 2024) and automated workflow design architectures (Hu et al., 2025; Zhang et al., 2024c; Zhou et al., 2025a).
@@ 2.4 Reasoning Regimes
Figure 6: Inference-time and training-time regimes of a reasoning system. We use tree search as an example to illustrate the inference scaling and trajectories collection. Given a query, inference scaling relies on extensive inference computation to improve the reasoner’s distribution. Specifically, it generates multiple candidate reasoning steps at each layer and selects the best solution to proceed (e.g., by using an external verifier or ensembling). In contrast, learning to reason focuses on collecting trajectories and training from the collected data with minimal inference-time computation. It takes all trajectories in the process (identical to those used in inference-scaling, allowing us to reuse the same tree) and labels them with preferences. The preference data can then be used to train the reasoner.
@@ 3.1.1 Constructing Reasoning Provoking Prompts
Although large-scale pre-training endows LLMs with patterns that support reasoning, these capabilities often remain latent under generic prompts. Liu et al. (2025c) demonstrate that deep-reasoning behaviors—such as reflection and self-verification, which signal profound analytical thought—can be amplified simply by increasing the sampling budget. This highlights the importance of designing prompts that deliberately provoke reasoning, thereby surfacing and leveraging the latent human priors within LLMs.
@@
Humans can address new problems by drawing analogy from relevant past experience (Holyoak, 2012). Inspired by this, Yasunaga et al. (2024) propose analogical prompting to guide LLMs to self-generate exemplars or knowledge relevant to the given problem as few-shot demonstrations for reasoning, outperforming hand-crafted or retrieved examples. For example, LLMs are prompted to generate a problem on calculating a third-order determinant before solving the given fourth-order determinant. Similarly, Chen et al. (2023d); Yang et al. (2023a); Luo et al. (2024a) highlight the effectiveness of self-generated relevant exemplars. Qin et al. (2025) further systematically assess the capability of LLMs to perform analogical reasoning and find that performance is not primarily determined by whether the exemplars are topically relevant to the task. Instead, they show that even exemplars from unrelated domains, such as self-generated biological exemplars, can lead to improved performance, as long as they are accurate and structurally aligned with the reasoning steps required by the target task. This highlights that the quality of the exemplar (its correctness, clarity, and structural usefulness for reasoning) can be the key limiting factor, rather than the relevancy regarding to the topic domain.
Conventionally, a fixed set of few-shot demonstrations is applied to all queries, which can be suboptimal, especially when queries vary significantly. An alternative approach is to retrieve demonstrations tailored to the current query. Research has shown that retrieval-based demonstration selection significantly improves task performance. The main goals for selecting demonstrations are similarity (Rubin et al., 2022; Agrawal et al., 2023; Li et al., 2023e; Ye et al., 2023a) and diversity (Levy et al., 2023; He et al., 2023; Kim et al., 2024a). Various retrieval strategies have been proposed for selecting $k$ demonstrations, including top-$k$ similarity-based retrieval (Liu et al., 2022; Li et al., 2023e), clustering-based retrieval (Luo et al., 2023c; Wang et al., 2024i), and iterative retrieval (Khattab et al., 2022; Levy et al., 2023; Wang et al., 2024e). These methods enable adaptive and effective demonstration selection, enhancing the model’s reasoning and generalization across diverse queries.
In addition, many-shot in-context learning has emerged as a complementary line of work, where hundreds or even thousands of demonstrations are provided to significantly enhance the performance of LLMs, especially on complex reasoning tasks (Li et al., 2023c; Agarwal et al., 2024; Zou et al., 2024; Gu et al., 2025). Many-shot prompting can be seen as an extreme form of demonstration engineering, where the focus is on scaling the quantity of demonstrations to maximize the model’s capacity to learn from in-context examples. However, the effectiveness of many-shot ICL is often limited by the high cost of obtaining a large number of labeled demonstrations. To mitigate this gap, Chen et al. (2025) recently introduce MAPLE, a novel influence-based many-shot ICL framework that identifies impactful unlabeled samples, pseudo-labels them by querying LLMs, and adaptively selects them for each test query. This approach effectively enhances many-shot ICL performance with minimal labeling cost, demonstrating improved adaptability and reasoning capabilities of LLMs.
@@ 3.1.2 Optimizing Reasoning Output with Search and Planning
Beyond subtasks, researchers emphasize the importance of robust reasoning structures such as hierarchical and decision-making processes that capture the underlying mechanisms involved in problem-solving. Zhou et al. (2024b) introduce Self-Discover, a framework that enables models to self-identify reasoning structures for any task using a seed set of general reasoning skill modules. Building on this, Aswani et al. (2024) propose Auto-Evolve, which dynamically adapts reasoning modules to accommodate more diverse problems. In addition to designing better reasoning steps, several studies address the need to correct intermediate steps. For example, Deng et al. (2024a); Yan et al. (2024) and Wu et al. (2024b) propose methods to refine intermediate outputs. Notably, Zhang et al. (2024i) observe that smaller models ($\leq$ 13B parameters) in particular need stronger models acting as verifiers to validate and correct intermediate steps.
@@ 3.2 Inference Scaling With Single-agent System
A pioneering approach in this domain is the ReAct framework (Yao et al., 2023b), which interleaves reasoning and acting by prompting LLMs to generate both reasoning traces and task-specific actions in an interleaved manner. This synergy allows the model to induce, track, and update action plans while interfacing with external sources (environment) to gather additional information. ReAct has demonstrated effectiveness across QA and interactive decision-making tasks. Building upon ReAct, LATS (Zhou et al., 2024a) unifies reasoning, acting, and planning within LLMs. By combining Monte Carlo Tree Search with ReAct, LATS enables structured search over a combinatorial space of reasoning and acting paths. More recently,  Liu et al. (2024f) formalize reasoning and acting with LLMs under a Bayesian adaptive MDP and propose RAFA, a theoretically grounded framework for orchestrating the reasoning and acting of LLMs.
@@ 3.2.2 Enhancement through Retrieval and Tool Utilization
Moreover, Anthropic introduced an open standard of Model Context Protocol (MCP) to seamlessly connect AI assistants with real-world data sources such as content repositories, business tools, and development environments. It provides a universal, scalable way for developers to create secure, two-way connections between AI tools and diverse data systems. While MCP holds significant promise, its adoption also introduces several challenges that must be addressed to support sustainable growth and responsible development. Hou et al. (2025) discussed some key issues, such as the absence of centralized security oversight, gaps in authentication and authorization, and difficulties in maintaining consistency across multi-step, cross-system workflows.
@@ 4.1.2 Reinforcement Learning for Reasoning
At each time step $t$, for a given state $s_{t}\in\mathcal{S}$, the agent selects an action $a_{t}$ and transitions to a new state $s_{t+1}$, receiving a reward $R(s_{t},a_{t})$ from the environment. The set of available actions at state $s_{t}$ may be restricted to a subset of $\mathcal{A}$, denoted $\mathcal{A}_{s_{t}}$ (i.e., $a_{t}\in\mathcal{A}_{s_{t}}$). In the context of autoregressive language modeling with LLMs, generally the next token depends on all the previous tokens. As such, in order to apply RL training for LLMs, one needs to define the states and actions of the problem such that they both satisfy the temporal dependency constraint of the language modeling task as well as the Markov property. One common approach is to define that the current state $s_{t}$ fully encapsulates all relevant information about the environment, in other words all previous tokens. This means the next state $s_{t+1}$ depends solely on the current state $s_{t}\in\mathcal{S}$ and the chosen action $a_{t}\in\mathcal{A}_{s_{t}}$. In this way, the current state no longer needs to retrieve information from the previous states to decide the next action. As such, the state transition is agnostic to the history or previous states and actions. Within this MDP framework, the goal of RL is to learn a policy model that selects optimal actions by maximizing the expected cumulative rewards (Eq. 1).
@@
As one of the primary variants of policy gradient methods, PPO has remained a popular and widely used RL algorithm (Schulman et al., 2017). To train the policy $\pi_{\theta}$, PPO utilizes two additional models: the reference model $\pi_{\theta_{\text{ref}}}$, which represents the initial state of the policy, and the value model $V$, which estimates the state value $V(s_{t})$. PPO begins by sampling a state-action trajectory $\tau$ with consecutive state-action pairs $s_{t+1}\sim(s_{t},a_{t})$, then collects the respective intermediate or process reward (if available) and final (outcome) reward. Then, it computes the advantage $A(s_{t},a_{t})$ of each action $a_{t}$ given the current state $s_{t}$, which is defined as the relative strength of that specific action $a_{t}$ compared to the probability-weighted actions that the policy could probably have taken from $s_{t}$. The advantage is formulated as
@@
where $Q(s_{t},a_{t})$ represents the expected cumulative total reward that the policy is expected to obtain if it takes action $a_{t}$ from $s_{t}$ and continue to follow the current policy, while $V(s_{t})$ denotes the expected total rewards obtainable from state $s_{t}$, known as the state value. The state value is equivalent to the expected value of $Q(s_{t},a^{\prime}_{t})$ marginalized over all probable actions the current policy $\pi_{\theta}$ may take from $s_{t}$. If $A(s_{t},a_{t})>0$, the action $a_{t}$ is encouraged, conversely, if $A(s_{t},a_{t})<0$, the action $a_{t}$ is discouraged. After computing the advantages, PPO optimizes the policy $\pi_{\theta}$ according to the following loss function.
@@
where $t\in[0,T]$ is a time step within trajectory $\tau$, $\pi_{\theta_{o}}$ is the fixed policy from previous episode or iteration, and $P$ is the transition distribution. The $clip$ function, applied to the probability ratio $\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{o}}(a_{t}|s_{t})}$, ensures that the policy does not deviate too drastically or rapidly from its previous version. This also help prevent catastrophic failure or suboptimal local solutions. Additionally, a KL divergence term $\mathcal{D_{\text{KL}}}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})$ is often incorporated into the loss function to constrain exploration during the later stages of training. $\pi_{\theta_{\text{ref}}}$ is often a fixed initial reference policy that we do not want our policy to deviate too much from, while $\pi_{\theta_{o}}$ is a snapshot of the current policy from the previous iteration which is updated regularly. Throughout the training process, both the policy $\pi_{\theta}$ and value model $V$ are iteratively updated.
@@
$L_{\text{REINFORCE}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}[(R(y,x)-b)\nabla_{\pi_{\theta}}\log\pi_{\theta}(y|x)]$ (5)
@@
$L_{\text{RLOO}}(\theta)=\frac{1}{k}\sum_{i=1}^{k}[R(y_{i},x)-\frac{1}{k-1}\sum_{j\neq i}R(y_{j},x)]\nabla_{\pi_{\theta}}\log\pi_{\theta}(y_{i}|x)$ (6)
where $k$ denotes the number of Monte Carlo samples. Unlike PPO, these algorithms do not rely on a parameterized value function (critic model) and instead depend solely on observed rewards. These methods share similarities with approaches such as Group-Relative Policy Optimization (GRPO) (Ramesh et al., 2024) and VinePPO (Kazemnejad et al., 2024), which will be discussed in detail below.
@@
Variants of GRPO, such as DAPO (Yu et al., 2025), have also been introduced to alleviate issues with GRPO like length bias and inappropriate penalties for responses that exceed the context length.